the length of a rectangle parking lot is 10 meters less than twice the width, the perimeter is 406 meters, find the length of the parking lot.
Answer by mananth(16949)

let the width be x
twice the width = 2x
10 less = 2x-10
the length is 2x-10
..
Perimeter of rectangle = 2L+2*W
2*x+2(2x-10)=406
2x+4x-20=406
6x=406+20
6x=426
x=71m which is the width.
length will be 2x-10
2*71-10
142-10
132m is the length.
..
CHECK
132+132+71+71=406 meters